OpenStudy (anonymous):

The vertical displacement, h(t),in metres, of a cork floating in rough water can be given by h(t)=0.5 cos(2t)+0.7sin(t), where t is the time , in seconds. Determine the maximum velocity that the cork achieves and the first time it occurs

OpenStudy (anonymous):

I found the first and second derivative, and then found the zeros of the second derivatives (so, where the velocity is highest) and then do i put those values back into the first derivative equation? Also, one of my zeros for the second derivative was negative so since I'm working with time I should discard that solution, correct?

OpenStudy (rational):

max/min values occur when first derivative = 0, right ?

OpenStudy (rational):

set \(h'(t) = 0\) and solve \(t\)

OpenStudy (rational):

that gives the critical points

OpenStudy (anonymous):

yes. but i'm looking for max and min velocity (first derivative is veolocity) so i'm looking for where the second derivative equals zero.

OpenStudy (rational):

Ahh I see, yes you're right !

OpenStudy (anonymous):

okay, I think I have everything right but I have an answers sheet and the answers are a little off so I just wanted to check that I was right.

OpenStudy (rational):

whats the maximum velocity you're getting ?

OpenStudy (anonymous):

my zeros for the second derivative are 0.927 and -0.675 (which i'm not using because negative time isn't a thing), so my max velocity is -0.53m/s

OpenStudy (anonymous):

which kind makes sense, because it's not like the cork can ever be out of the water while it's floating, so it can never have a positive value.

OpenStudy (rational):

im getting max velocity = +1.522, when t = 5.6

OpenStudy (anonymous):

my teacher had 5.6 as an asnswer too! can you tell me how you got 5.6?

OpenStudy (rational):

whats the period of velocity ?

OpenStudy (rational):

velocity = -sin(2t) + 0.7cos(t) period = pi + 2pi = 3pi, right ?

OpenStudy (rational):

so you need to look for ALL critical points in the interval : t = [0, 3pi]

OpenStudy (rational):

and pick the max value

OpenStudy (rational):

your -0.53 is just a local minimum value.

OpenStudy (rational):

To get the Global maximum value : 1) Find all the critical values 2) Take max of above.

OpenStudy (rational):

since the period of velocity is 3pi, finding all the critical values in [0, 3pi] is sufficient